OpenStudy (anonymous):

write y=x-4 in standerd form?

OpenStudy (unklerhaukus):

if by standard form you mean \[y=mx+b\]your equation is already in standard form gradient \[m=1\] y intercept\[b=-4\]

OpenStudy (anonymous):

x -y = 4

OpenStudy (mertsj):

Standard form is ax + by = c

OpenStudy (anonymous):

yea...sure its a standard form of straight line in form of y=mx+c with slope "m" & intercept as"c"

OpenStudy (anonymous):

Not sure if it's a difference due to countries, but we (Australians) refer to GENERAL form, which is ax + by + c = 0

OpenStudy (anonymous):

no its a equation for family of st line ...............

OpenStudy (unklerhaukus):

@steppnstrong but i am australian

OpenStudy (anonymous):

i believe y = mx + b is generally reffered to as slope-intercept form, due to the fact that it gives you the slope and the y-intercept a general/standard form would be ax + by = c, or ax + bc - c = 0. , as it gives you two arbitrary points on the line (actually intercepts) and allows you to solve for other forms. so the answer is: y=x-4 <--slope intercept y - x = -4 <-- subtract x from both sides and get the general form (answer)

Directrix (directrix):

1 x + (-1) y = 4 Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html